### v1

# Construction of Functional Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model

Abstract
The convergence of materials science and artificial intelligence has unlocked new opportunities for gathering, analyzing, and generating novel materials sourced from extensive scientific literature. Despite the potential benefits, persistent challenges such as manual annotation, precise extraction, and traceability issues remain. Large language models have emerged as promising solutions to address these obstacles. This paper introduces Functional Materials Knowledge Graph (FMKG), a multidisciplinary materials science knowledge graph. Through the utilization of advanced natural language processing techniques, extracting millions of entities to form triples from a corpus comprising all high-quality research papers published in the last decade. It organizes unstructured information into nine distinct labels, covering Name, Formula, Acronym, Structure/Phase, Properties, Descriptor, Synthesis, Characterization Method, Application, and Domain, seamlessly integrating papers’ Digital Object Identifiers. As the latest structured database for functional materials, FMKG acts as a powerful catalyst for expediting the development of functional materials and a fundation for building a more comprehensive material knowledge graph using full paper text. Furthermore, our research lays the groundwork for practical text-mining-based knowledge management systems, not only in intricate materials systems but also applicable to other specialized domains.
Introduction
In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions. These conventional research approaches significantly impede the life-cycle of high-performance material research. Given the specialization, inherent complexity, and vast knowledge base of material science, researchers focusing on a single direction often struggle to efficiently access and understand material knowledge from multidisciplinary studies. For instance, researchers in solar cell development might not fully comprehend studies related to solid-state batteries or organic light-emitting diodes. Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other. To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [1]. However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering. This fragmentation of knowledge represents a significant barrier to interdisciplinary collaboration and innovation. A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information of material science from the vast ocean of literature remains demanding. To obtain clearer sense of materials properties some structured database project such as Materials Project [2], OQMD [3], and NOMAD [4] were developed. However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [5]. While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations. Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
Furthermore, the complexity of materials information extends beyond composition and structure to encompass their respective application fields. For instance, organic materials are commonly utilized in biological applications, semiconductors serve as integral components in electronics, and metals find applications in mechanical engineering. The process of designing novel materials typically commences with a clear understanding of their intended applications, aiming to maximize research efficiency. Given this context, compared to universal databases like the Materials Project, which might not be as useful due to their broad focus, specific databases focused on applications or traits hold the potential to provide more valuable information for researchers in relevant industries. This underscores the importance of developing dedicated databases that cater to the nuanced needs of materials science research, facilitating a more targeted approach to material discovery and application.
Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide. The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [6], [7]. These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [8]. However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [9]. This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [10]. With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [11]. For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [12]. Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [13]. A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [14]. Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [1], [15].
However, these material knowledge graphs face even greater challenges. Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [16]. Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [17]. This can diminish the authenticity and credibility of the KG. Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science. Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the extraction and credibility of structured information [18], [19]. The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [20],[21]. This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
In this paper, we have achieved significant advancements in the development of Functional Materials Knowledge Graph (FMKG), a pioneering graph database tailored for the field of functional materials. Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy. Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet. This method also makes updating KG very convenient. 2) We constructed the first knowledge graph dedicated for functional materials, where researchers can easily get the information about functional material through query the FMKG. 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs.
Methods
Figure 1(a) illustrates the comprehensive workflow of our research. Through NER and RE tasks, we extract the structural information about catalyst, battery and solar cell. After ER and normalization, we integrate information from these three fields and construct a knowledge graph. Specifically, the pipeline displayed in the Figure 1(b), commencing with the manual annotation and normalization of the initial training data set, prompting the fine-tuned LLMs specifically for NER and RE tasks. This inference dataset is subsequently divided into ten batches, a crucial step for the iterative process that follows. Then, we finish ER task through the NLP technology including ChemDataExactor [22], mat2vec [23] and our expert dictionary. After ER, high-quality results are meticulously selected to augment the training set, thereby enhancing the model’s performance in subsequent iterations. Finally, the knowledge graph is constructed using the triples transferred from normalizd result after the last iteration.
Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
Data preparation and schema design
Material experts undertook the annotation of nine distinct categories from 75 abstracts of research papers, act as LLM’s training dataset. These categories encompassed core labels - "Name," "Formula," or "Acronym" - which consistently represent specific materials across diverse papers, albeit in varied forms. Additionally, supplementary labels included "Descriptor," "Structure/Phase," "Application," "Property," "Synthesis," and "Characterization." The "Property" label served to delineate quantitative properties such as ’length,’ ’specific surface area,’ and ’mass,’ whereas the "Descriptor" label encapsulated qualitative attributes like ’stable,’ ’vertically,’ and ’safe.’ Since from the application aspects, developers always concern about its own properties, we also annotated the essential parameters as "Property" directly linked to specific "Application". For instance, attributes like ’specific capacity’ for lithium-ion batteries, ’energy conversion efficiency’ for solar cells, and ’H2 production rate’ for hydrogen evolution reaction were systematically annotated.
As the Figure 2 shows, the central node, "Material," is pivotal and connects to nodes that describe its nomenclature, composition, and various attributes. Specifically, "Material" is consisted of "Formula", "Name" and "Acronym". "Structure/Phase" describing its physical form; "Application" denoting its practical uses; and "Property" outlining its inherent characteristics. The "Material" node is also associated with "Descriptor" node that provide additional qualitative information. Moreover, the "Application" node branches out to "Property", "Descriptor" and a "Domain" suggesting further specification and contextual relevance of application. To save the source of nodes and relations, each node is connected to a "Digital Object Identifier (DOI)" node. If we want to determine the source article from which the relation is derived, we can obtain the result by querying the intersection of "DOI" node neighbors at both ends of the relation.
Figure 2: This schematic represents the FMKG schema (a) All the nodes in FMKG connect to all their sources - "DOI". (b) Relations between all labels. (c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction. This normalization ensured uniform representation of entities with similar meanings. For example, terms such as ’Lithium-Ion Battery’ and ’Li-ion batteries’ were standardized to ’lithium-ion battery,’ while phrases like ’solution casting method,’ ’solvent post-treatment method,’ and ’solution-based deposition’ were simplified to ’solution-processed.’ All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs. The field of functional materials encompasses a wide range of areas. At this stage, our priority is to focus on energy materials. We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science. Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
LLMs training, evaluation and inference
The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA7b, LLaMA2 7b, and Darwin [24]. Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data. The models underwent training over 10 epochs with a batch size of 1. Additionally, 60 abstracts were annotated to assess the LLMs" performance. Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets. Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities. Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
Precision=True Positives (tp)True Positives (tp) + False Positives ($\text{Precision}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) +False Positives (fp)}}$ (1)
Recall=True Positives (tp)True Positives (tp) + False Negatives ($\text{Recall}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) + False Negatives (fn)}}$ (2)
$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
For NER, a true positive (tp) is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema. Conversely, a false positive (fp) occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative (fn) is when the model fails to identify a correct entity, relation, or schema element that should have been recognized. After definition, we can evaluate each task using Equations (1), (2), and (3).
After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts. The output of inferences are organized in the "DOI - text - response" format. Consequently, the fine-tuned LLM not only extracts entities but also assigns them to appropriate labels, thereby accomplishing NER and RE tasks concurrently. Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
Entity resolution
The quality of KG is crucial for its credibility, we need to further check and correct the inference results before graph construction. For the core label, we divide standardization into two steps: ER-NF/A to ER-N/F. In stage of ER-NF/A, we apply the ChemDataExtractor to derive the Formula and pairs of "Name - Acronym" from the abstract. Subsequently, we utilize the mat2vec model to perform word embedding on entities extracted by both the LLM and ChemDataExtractor. Through comparing their similarities, we aim to correct the core entities to guarantee their precision and the accurate correspondence between "Name" and "Acronym". However, The Name and Formula of materials are often placed in incorrect labels and are difficult to distinguish through ChemDataExtractor. Therefore, we conduct ER-N/F. We selected 2000 correctly classified labels as the training set and used to fine-tune LLM to complete a binary classification problem. Additionally, we selected other 200 labels to evaluate the performance of binary classification.
To normalize entities and relations in other label (ER-OL), we conduct a Word Embedding Clustering to create a expert dictionary. We convert words into vector form using mat2vec and group semantically similar words into the same cluster based on their vector similarity. Specifically, we develop a Density-Based Dynamic Vector Clustering method that dynamically forms clusters based on the similarity of vector representations, without predefining the number of clusters. This approach utilizes a similarity threshold to determine cluster membership, allowing entities to join existing clusters or form new ones based on their proximity in the feature space, effectively adapting to the inherent structure of the data.
By naming each cluster through experts in the field of material science, we extracted a total of approximately 600 vocabulary, including the structure, phase, applications, characterization methods, synthesis methods, properties, and descriptions of most energy materials. Subsequently, we adopt analytic strategies and embed the vocabulary from expert dictionary and the enetities extracted by LLM using mat2vec. After similarity comparison, we finish the standardization and ensure that both entities and relations are correct. It is worth mentioning that the content of "Property" and "Descriptor" is often not limited to the energy field and may cover more information. Therefore, we implement relatively tolerant standardization for these two labels, which only normalize them with matching content but does not delete unmatched entities and relations.
To improve the caliber of the training dataset, we select a small number of high-quality data from the normalized inference output of each iterative process, incorporating this data into the training set. High-quality data is characterized by its high precision and recall metrics. Furthermore, the performance of the fine-tuned LLM, trained on each progressively augmented training set, is rigorously evaluated to ensure optimal effectiveness.
Knowledge graph construction
To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head, the names of other labels forming the relations, and their values acting as tails. Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym". Furthermore, each head and tail node is linked to the DOI of the source article associated with the triplet. This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities. Then we transfer the triples into FMKG and store the FMKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions. To facilitate access to the detailed information, we also make the dataset available in the CSV format for straightforward data handling.
Result
In Table 1, the performance comparison of LLMs includes Darwin, LLaMA, and LLaMA2 in NER, RE and ER task within the materials science domain is presented. Darwin achieves markedly higher F1 scores in both NER and RE tasks than the LLaMA 7b and LLaMA2 7b models. This indicates that Darwin can perform more satisfactory results in text tasks in the field of materials science. However, there has been no significant difference in their effectiveness for ER tasks between these models. We speculate that this is due to insufficient contextual memory capabilities of the LLMs.
In response to the challenge, entity resolution was applied to the inference outputs, which substantially improved the performance in Standardization (ER) tasks. This refinement process has also led to enhancements in the NER and Relation RE tasks as well. The improvements can be attributed to the ER task removing some of the incorrect inferences that resulted from hallucinations and providing a secondary adjustment of the extracted relations.
Model Task Precision Recall F1 score
NER 0.6101 0.6216 0.6158
Llama 7b RE 0.5305 0.5405 0.5355
ER 0.3687 0.3757 0.3722
NER 0.7419 0.7667 0.7541
Llama2 7b RE 0.6452 0.6667 0.6557
ER 0.4484 0.4633 0.4557
NER 0.8013 0.7935 0.7974
Darwin RE 0.7036 0.6968 0.7002
ER 0.4593 0.4548 0.4571
NER 0.9520 0.9083 0.9296
Darwin (Normalized) RE 0.9039 0.8625 0.8827
ER 0.9127 0.8708 0.8913
Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
To gain a more comprehensive understanding of the improvements brought by each step in the standardization process, we conducted ablation experiments, the results are shown in Table 2 which clearly shows that each method contributes to the improvement of ER. The result indicates the most effective method is to optimize through the expert dictionary, which indicates that in addition to the accuracy of the expert dictionary, it can also fully cover all other labels except for the three labels representing the material. The improvement of ER-NF/A is also significant, as this process can eliminate most erroneous materials, and even has a stronger improvement on NER than ER-OL.
  Method NER F1 $\nabla$ RE F1 $\nabla$ ER F1 $\nabla$
Darwin (Normalized) 92.96 - 88.27 - 89.13 -
ER-N/F 92.96 - 83.29 -4.98 89.13 -
ER-NF/A 84.01 -8.85 83.07 -5.20 84.54 -4.59
ER-OL 88.59 -4.37 80.20 -8.07 50.00 -38.83
Table 2: Result of the ablation experiment in normalization.
Besides, ER-N/F pay attention to adjust the relation error between "Name" and "Formula", from a contribution perspective, it is also an indispensable part of ER. We also evaluated the LLM classifier by selecting 200 manually labeled data, and the statistical evaluation results in Table 3 shows the strong performance of LLM used in classification task.
Model Accuracy F1 Score Specificity
Darwin 0.9650 0.9581 0.9741
Table 3: Binary classification performance through Fine-tuned Darwin.
The triples are transferred from normalized data to construct the FMKG which contains 162,605 nodes and 731,772 edges, the schematic diagram shows in Figure 3. In the graph, there are 11 types of nodes and 13 types of relations. The FMKG store large amount of information in functional materials includes battery, catalyst and solar cell. For example, as the Figure 4 shows, by analyzing the frequency of material appearances in the battery domain, we observe that Co2O3 is the most frequently occurring material, followed by MoS2, graphite, TiO2, LiCoO2 etc which are commonly used in the battery domain. Likewise, it is evident that the most prevalent application within the battery field is lithium-ion batteries, which appear with a frequency significantly higher than other applications in the battery domain. These insights align with our understanding of these domains and materials, confirming that our knowledge graph effectively stores factual knowledge and offers a knowledge platform for material researchers.
Figure 3: (a) Global schematic diagram of FMKG; (b) Local schematic diagram of FMKG.
Figure 4: The image shows the top ten materials and applications in each field.
For further validation of FMKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science. After splitting these triples apart, annotator can obtain 1000 entities and 500 relations. The report from annotator shows in Table 4.
“Name” is the highest priority core label and only serves as the head in the knowledge graph, so it will not be used as a relation. Notably, the labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation. This is due to the rigorous standardization of entities under these categories, using our expert dictionary. In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation. But the results are still satisfactory.
"Name" and "Acronym" has also achieved a commendable 100% entity accuracy. However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations. A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor. The impact of this kind of error is not significant, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
Label Number in total Entity disagree % disagree Relation disagree % disagree
Formula 257 0 0 N/A N/A
Name 227 0 0 12 5.3
Acronym 26 0 0 3 11.5
Descriptor 127 7 5.5 12 9.4
Property 147 12 8.2 9 6.1
Application 73 0 0 0 0
Structure/Phase 36 0 0 0 0
Synthesis 50 0 0 0 0
Characterization 57 0 0 0 0
Table 4: Human evaluation metric on randomly selected triples.
Discussion
In this research, we construct a functional material KG by fine-tuned LLMs and ensuring the traceability of all information. During the process, we encountered two main challenges and proposed corresponding solution: 1) Fine tuning LLM reduces the requirement for training set volume, but requires higher quality for each data in training set. when the quality of training dataset is limited, the performance of NER and RE is not good. Recursively generating training set data automatically from inference result is a good solution. High recall and precision data from each inference can be put into the training set after simple manual check. 2) Fine-tuned LLMs perform well on NER and RE task, but for ER task, they can not shows a same. However, the ER is also important in graph construction. To address the issue, we apply several ER methods and technologies including ChemDataExtractor, mat2Vec, Word Embedding Clustering and construct an expert dictionary storing information about energy material. Theses methods and technologies can not only achieve high perfromance ER task, but also remove some erroneous entities and relations.
To better demonstrate the reliability of our method and FMKG, we compare them with existing structural information extraction method and material KG. We employ a similar approach to NERRE task of John Dagdelen et al. [21] to extract structured information from the text. The difference is that we have applied ER method to correct entities and relations, and finish ER tasks before graph construction. We evaluated our pipeline on NERRE’s evaluation dataset. It worth to mention that cross dataset evaluation allow for flexibility of multiple plausible annotations of an entity under different labels (e.g., "cathode" in NERRE is placed in "Application" label while our expert tend to place it in "Descriptor" since "cathode" and "anode" are component of "battery" which is an application.), and presented the results in a Table 5. From the results, it is evident that our pipeline has slightly lower recall across labels compared to the NERRE task, but our method holds a competitive edge in precision. This is because we believe that the primary goal of KG construction is to ensure the accuracy of the data. Hence, in the process of ER, our strategy tends to sacrifice a portion of recall in exchange for higher precision. However, when considering both recall and precision, our method scores higher in F1 across the "Acronym", "Application", "Structure/Phase", and "Descriptor" labels compared to NERRE, with the improvement in "Acronym" being particularly significant due to its lower precision in NERRE. This underscores the advantage of our ER pipeline and indirectly attests to the credibility of the FMKG.
Label Extraction recall (NERRE) Extraction precision (NERRE) Extraction F1 (NERRE) Extraction recall (DarwinN) Extraction precision (DarwinN) Extraction F1 (DarwinN)
Formula 0.943 0.943 0.943* 0.828 1.000 0.906
Name 0.692 1.000 0.818 0.667 1.000 0.800
Acronym 0.667 0.400 0.500 0.571 1.000 0.727
Descriptor 0.576 0.905 0.704 0.781 0.962 0.862
Application 0.797 0.870 0.832 0.750 1.000 0.857
Structure/Phase 0.754 0.920 0.829 0.613 0.905 0.730
*Best F1 scores for each label.
Table 5: Evaluation comparison of NERRE and DarwinN.
Knowledge graphs specifically dedicated to general materials are rare. In our discussion, we draw a comparison with MatKG2 [15], showcasing the differences in construction methodologies between MatKG2 and FMKG, as depicted in the Figure 5. The construction of MatKG2 involves in a multi-step NER and RE pipeline, which can not catch the source of relations. In our method, we use end-to-end method that use a single LLM have been investigated for joint named entity recognition and relation extraction. It can not only catch keep source of each triple in KG so that enhance the factual basis of KG, but also perform better on RE task.
Figure 5: Schematic comparison of FMKG and MatKG2.
KGs are always dynamic, especially in cutting-edge field such as material science. The LLM internalizes a large amount of knowledge during the training process, it can perform good inference even when new nodes and relationships appear, making KG easy to update. Besides, our work is a starting point where other material researchers can continue build FMKG using our result as training dataset, using our dictionary, ChemDataExtractor and mat2vec to achieve ER task, our method can also be extend to full paper, extract more functional material information to construct a larger and more comprehensive functional material knowledge graph. Besides, our pipeline can also be used in other specialized domains to construct domain-specific KGs.
Conclusion
In this study, we introduce a new NLP pipeline for KG construction, which aim to efficiently extract the triples from unstructured scientific texts. The main feature of the method is that it can fine-tune the LLMs by annotating a small amount of data, and use the fine-tuned LLM to extract structured information from a large amount of unstructured text. The entire process does not rely on any prediction, which can maximize the authenticity and traceability of structured information. By employing this method, we construct a Functional Material Knowledge Graph (FMKG) contains the materials and their related knowledge from abstract of 150,000 peer-reviewed paper. After analyzing, we have demonstrated the effectiveness and credibility of FMKG.
In addition, our method and KG have great potential in different dimensions. Firstly, enhancing the depth of structured information extraction to encompass entire research papers promises a richer, more detailed knowledge graph. This involves not only expanding the scope of data analyzed but refining the process to capture nuances within complex scientific texts. Secondly, refining entity labels within our system allows for a more precise categorization of data, including incorporating detailed attributes such as synthesis conditions or property parameters, which significantly improves the granularity and utility of the knowledge graph. Thirdly, the versatility of our NLP pipeline suggests its applicability across different scientific domains, offering a template for constructing domain-specific knowledge graphs beyond material science. Lastly, integrating FMKG with existing knowledge graphs like MatKG opens avenues for creating a more interconnected and comprehensive dataset, facilitating advanced research and application development in material science and beyond.

### v2

@@
-Construction of Functional Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model
+Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model
 Abstract
-The convergence of materials science and artificial intelligence has unlocked new opportunities for gathering, analyzing, and generating novel materials sourced from extensive scientific literature.
-Despite the potential benefits, persistent challenges such as manual annotation, precise extraction, and traceability issues remain.
-Large language models have emerged as promising solutions to address these obstacles.
-This paper introduces Functional Materials Knowledge Graph (FMKG), a multidisciplinary materials science knowledge graph.
-Through the utilization of advanced natural language processing techniques, extracting millions of entities to form triples from a corpus comprising all high-quality research papers published in the last decade.
-It organizes unstructured information into nine distinct labels, covering Name, Formula, Acronym, Structure/Phase, Properties, Descriptor, Synthesis, Characterization Method, Application, and Domain, seamlessly integrating papers’ Digital Object Identifiers.
-As the latest structured database for functional materials, FMKG acts as a powerful catalyst for expediting the development of functional materials and a fundation for building a more comprehensive material knowledge graph using full paper text.
-Furthermore, our research lays the groundwork for practical text-mining-based knowledge management systems, not only in intricate materials systems but also applicable to other specialized domains.
+Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges for efficient discovery and integration of new materials.
+Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation.
+Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information.
+To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques, integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges.
+MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration.
+By implementing network-based algorithms, MKG not only facilitates efficient link prediction but also significantly reduces reliance on traditional experimental methods.
+This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
 Introduction
@@
 Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
-Furthermore, the complexity of materials information extends beyond composition and structure to encompass their respective application fields.
-For instance, organic materials are commonly utilized in biological applications, semiconductors serve as integral components in electronics, and metals find applications in mechanical engineering.
-The process of designing novel materials typically commences with a clear understanding of their intended applications, aiming to maximize research efficiency.
-Given this context, compared to universal databases like the Materials Project, which might not be as useful due to their broad focus, specific databases focused on applications or traits hold the potential to provide more valuable information for researchers in relevant industries.
-This underscores the importance of developing dedicated databases that cater to the nuanced needs of materials science research, facilitating a more targeted approach to material discovery and application.
 Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide.
@@
 This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
-In this paper, we have achieved significant advancements in the development of Functional Materials Knowledge Graph (FMKG), a pioneering graph database tailored for the field of functional materials.
+In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science.
 Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy.
@@
 This method also makes updating KG very convenient.
-2) We constructed the first knowledge graph dedicated for functional materials, where researchers can easily get the information about functional material through query the FMKG.
+2) We constructed the first accurate knowledge graph dedicated for materials, where researchers can easily get the information about material through query the MKG.
 3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs.
+4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
 Methods
-Figure 1(a) illustrates the comprehensive workflow of our research.
-Through NER and RE tasks, we extract the structural information about catalyst, battery and solar cell.
-After ER and normalization, we integrate information from these three fields and construct a knowledge graph.
-Specifically, the pipeline displayed in the Figure 1(b), commencing with the manual annotation and normalization of the initial training data set, prompting the fine-tuned LLMs specifically for NER and RE tasks.
-This inference dataset is subsequently divided into ten batches, a crucial step for the iterative process that follows.
-Then, we finish ER task through the NLP technology including ChemDataExactor [22], mat2vec [23] and our expert dictionary.
-After ER, high-quality results are meticulously selected to augment the training set, thereby enhancing the model’s performance in subsequent iterations.
-Finally, the knowledge graph is constructed using the triples transferred from normalizd result after the last iteration.
+Figure 1(a) presents the elaborate workflow of our study.
+Initially, through Named Entity Recognition (NER) and Relation Extraction (RE) tasks, we extract structural information pertinent to catalysts, batteries, and solar cells.
+Subsequent Entity Resolution (ER) and data normalization steps enable the integration of information from these distinct fields into a unified knowledge graph.
+Specifically, as outlined in Figure 1(b), the process begins with the manual annotation and normalization of the initial training dataset, essential for fine-tuning Large Language Models (LLMs) tailored for NERRE tasks.
+The inference dataset is then methodically divided into ten batches, facilitating the iterative refinement process that follows.
+The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [22], mat2vec [23], and our expertly curated dictionary.
+After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations.
+The knowledge graph is finally constructed using the normalized data from the last iteration.
+To complete the graph and predict potential material applications, we employ both network-based algorithms and graph embeddings.
+This methodology provides critical insights and recommendations for researchers in the materials science domain.
 Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
 Data preparation and schema design
-Material experts undertook the annotation of nine distinct categories from 75 abstracts of research papers, act as LLM’s training dataset.
-These categories encompassed core labels - "Name," "Formula," or "Acronym" - which consistently represent specific materials across diverse papers, albeit in varied forms.
-Additionally, supplementary labels included "Descriptor," "Structure/Phase," "Application," "Property," "Synthesis," and "Characterization." The "Property" label served to delineate quantitative properties such as ’length,’ ’specific surface area,’ and ’mass,’ whereas the "Descriptor" label encapsulated qualitative attributes like ’stable,’ ’vertically,’ and ’safe.’ Since from the application aspects, developers always concern about its own properties, we also annotated the essential parameters as "Property" directly linked to specific "Application".
-For instance, attributes like ’specific capacity’ for lithium-ion batteries, ’energy conversion efficiency’ for solar cells, and ’H2 production rate’ for hydrogen evolution reaction were systematically annotated.
-As the Figure 2 shows, the central node, "Material," is pivotal and connects to nodes that describe its nomenclature, composition, and various attributes.
-Specifically, "Material" is consisted of "Formula", "Name" and "Acronym".
-"Structure/Phase" describing its physical form; "Application" denoting its practical uses; and "Property" outlining its inherent characteristics.
-The "Material" node is also associated with "Descriptor" node that provide additional qualitative information.
-Moreover, the "Application" node branches out to "Property", "Descriptor" and a "Domain" suggesting further specification and contextual relevance of application.
-To save the source of nodes and relations, each node is connected to a "Digital Object Identifier (DOI)" node.
-If we want to determine the source article from which the relation is derived, we can obtain the result by querying the intersection of "DOI" node neighbors at both ends of the relation.
-Figure 2: This schematic represents the FMKG schema (a) All the nodes in FMKG connect to all their sources - "DOI".
+Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs).
+The structure of the data reflects the structure of the KG.
+As illustrated in Figure 2, the "Material" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes.
+Specifically, the "Material" node comprises sub-nodes such as "Formula," "Name," and "Acronym" for identifying materials.
+The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics.
+Additionally, the "Descriptor" node provides qualitative information enhancing the contextual understanding of the material.
+Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application.
+To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node.
+This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations was derived, ensuring data integrity within the graph structure.
+Figure 2: This schema represents the ontology of MKG (a) All the nodes in MKG connect to all their sources - "DOI".
 (b) Relations between all labels.
@@
 This normalization ensured uniform representation of entities with similar meanings.
-For example, terms such as ’Lithium-Ion Battery’ and ’Li-ion batteries’ were standardized to ’lithium-ion battery,’ while phrases like ’solution casting method,’ ’solvent post-treatment method,’ and ’solution-based deposition’ were simplified to ’solution-processed.’ All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs.
-The field of functional materials encompasses a wide range of areas.
+For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed", ensuring consistency and facilitating effective training of LLMs.
+The field of materials encompasses a wide range of areas.
 At this stage, our priority is to focus on energy materials.
@@
 LLMs training, evaluation and inference
-The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA7b, LLaMA2 7b, and Darwin [24].
+The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [24], and Darwin [25].
 Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data.
@@
 Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
-Precision=True Positives (tp)True Positives (tp) + False Positives ($\text{Precision}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) +False Positives (fp)}}$ (1)
-Recall=True Positives (tp)True Positives (tp) + False Negatives ($\text{Recall}=\frac{\text{True Positives (tp)}}{\text{True Positives (tp) + False Negatives (fn)}}$ (2)
-$F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
-For NER, a true positive (tp) is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema.
-Conversely, a false positive (fp) occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative (fn) is when the model fails to identify a correct entity, relation, or schema element that should have been recognized.
-After definition, we can evaluate each task using Equations (1), (2), and (3).
+For NER, a true positive is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema.
+Conversely, a false positive occurs when the model incorrectly identifies an entity, relation, or standardization, and a false negative is when the model fails to identify a correct entity, relation, or schema element that should have been recognized.
+After defining these terms, we can evaluate each task using standard precision, recall, and F1 score metrics.
 After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts.
@@
 Entity resolution
-The quality of KG is crucial for its credibility, we need to further check and correct the inference results before graph construction.
-For the core label, we divide standardization into two steps: ER-NF/A to ER-N/F.
-In stage of ER-NF/A, we apply the ChemDataExtractor to derive the Formula and pairs of "Name - Acronym" from the abstract.
-Subsequently, we utilize the mat2vec model to perform word embedding on entities extracted by both the LLM and ChemDataExtractor.
-Through comparing their similarities, we aim to correct the core entities to guarantee their precision and the accurate correspondence between "Name" and "Acronym".
-However, The Name and Formula of materials are often placed in incorrect labels and are difficult to distinguish through ChemDataExtractor.
-Therefore, we conduct ER-N/F.
-We selected 2000 correctly classified labels as the training set and used to fine-tune LLM to complete a binary classification problem.
-Additionally, we selected other 200 labels to evaluate the performance of binary classification.
-To normalize entities and relations in other label (ER-OL), we conduct a Word Embedding Clustering to create a expert dictionary.
-We convert words into vector form using mat2vec and group semantically similar words into the same cluster based on their vector similarity.
-Specifically, we develop a Density-Based Dynamic Vector Clustering method that dynamically forms clusters based on the similarity of vector representations, without predefining the number of clusters.
-This approach utilizes a similarity threshold to determine cluster membership, allowing entities to join existing clusters or form new ones based on their proximity in the feature space, effectively adapting to the inherent structure of the data.
-By naming each cluster through experts in the field of material science, we extracted a total of approximately 600 vocabulary, including the structure, phase, applications, characterization methods, synthesis methods, properties, and descriptions of most energy materials.
-Subsequently, we adopt analytic strategies and embed the vocabulary from expert dictionary and the enetities extracted by LLM using mat2vec.
-After similarity comparison, we finish the standardization and ensure that both entities and relations are correct.
-It is worth mentioning that the content of "Property" and "Descriptor" is often not limited to the energy field and may cover more information.
-Therefore, we implement relatively tolerant standardization for these two labels, which only normalize them with matching content but does not delete unmatched entities and relations.
-To improve the caliber of the training dataset, we select a small number of high-quality data from the normalized inference output of each iterative process, incorporating this data into the training set.
-High-quality data is characterized by its high precision and recall metrics.
-Furthermore, the performance of the fine-tuned LLM, trained on each progressively augmented training set, is rigorously evaluated to ensure optimal effectiveness.
+The quality of KG is crucial for its credibility, to further check and correct the inference results before graph construction.
+To ensure the precision of these results, we initially employ ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts.
+Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model.
+We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations.
+Given the frequent mislabeling of "Name" and "Formula", we progress to the Entity Resolution - Name/Formulas (ER-N/F) phase.
+Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
+For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [26] algorithm to create the expert dictionary (ER-ED).
+This algorithm dynamically forms clusters based on vector similarity, without the need to predefine the number of clusters.
+Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more.
+This dictionary, along with the entities extracted by the LLM, undergoes similarity analysis to standardize and confirm the accuracy of both entities and relations.
+To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
 Knowledge graph construction
-To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head, the names of other labels forming the relations, and their values acting as tails.
+Given a collection of entities $E$ and relations $R$, a knowledge graph $\mathcal{K}=E\times R\times E$ is structured as a directed multi-relational graph.
+It comprises triples formatted as $(h,r,t)\in\mathcal{K}$, where $h$ and $t$ are entities within $E$.
+To structure the inference results into triples, we employ three labels that signify the material as the core label, with the core label serving as the head $h$, the names of other labels forming the relations $R$, and their values acting as tails $t$.
 Within the hierarchy of core labels, "Formula" is accorded the highest priority, followed by "Name", and then "Acronym".
-Furthermore, each head and tail node is linked to the DOI of the source article associated with the triplet.
+Furthermore, each head $h$ and tail $t$ node is linked to the DOI of the source article associated with the triplet.
 This setup allows us to ascertain the provenance of the relation between any two nodes by examining the intersection in their connected entities.
-Then we transfer the triples into FMKG and store the FMKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions.
-To facilitate access to the detailed information, we also make the dataset available in the CSV format for straightforward data handling.
+Then we transfer the triples into MKG and store the MKG via graph database Neo4j, which also supports the subgraph matching function, where subgraph matching naturally suits the need to search for certain materials with user-input conditions.
+To facilitate access to the detailed information, we also make the dataset available in the RDF and CSV format for straightforward data handling.
+Graph completion
+In the analysis depicted in Figure 3, we enhance the traditional Jaccard similarity, hereafter referred to as $J$, to form a comprehensive similarity metric for evaluating potential material-to-application links.
+The Jaccard similarity is mathematically defined as:
+$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
+where $A$ and $B$ represent the attribute sets of two compared materials, such as their properties or applications.
+To refine our predication method, we extend this basic formulation by incorporating weighted factors that reflect the relevance and impact of each attribute, resulting in a modified similarity score for "Materials-Applications", $J_{\text{mat}}$.
+This score is computed as follows:
+$J_{\text{mat}}(A,B)=\sum_{i=1}^{n}w_{i}\cdot\frac{|A_{i}\cap B_{i}|}{|A_{i}\cup B_{i}|}$
+where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application.
+This approach mimics the investigative processes commonly employed by materials scientists when developing new materials.
+In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field.
+This enhanced method not only accounts for the quantity of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application.
+A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application.
+Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
+Figure 3: The process of network-based graph completion.
+In addition to the Enhanced Jaccard Similarity, we employ the TransE model [27].
+TransE treats entities and relations as vectors in the same embedding space.
+The core idea of TransE is to model relations by interpreting them as translations in the embedding space.
+For a given relation $r$ between head $h$ and tail $t$, the relationship is said to hold if:
+$\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|\approx 0$
+[…]
 Result
-In Table 1, the performance comparison of LLMs includes Darwin, LLaMA, and LLaMA2 in NER, RE and ER task within the materials science domain is presented.
-Darwin achieves markedly higher F1 scores in both NER and RE tasks than the LLaMA 7b and LLaMA2 7b models.
-This indicates that Darwin can perform more satisfactory results in text tasks in the field of materials science.
-However, there has been no significant difference in their effectiveness for ER tasks between these models.
-We speculate that this is due to insufficient contextual memory capabilities of the LLMs.
-In response to the challenge, entity resolution was applied to the inference outputs, which substantially improved the performance in Standardization (ER) tasks.
-This refinement process has also led to enhancements in the NER and Relation RE tasks as well.
-The improvements can be attributed to the ER task removing some of the incorrect inferences that resulted from hallucinations and providing a secondary adjustment of the extracted relations.
+Figure 4: Schematic comparison of MKG and MatKG2.
+To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [15], highlighting the differences in their construction methodologies as illustrated in Figure 4.
+MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation.
+In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously.
+This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
+To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in the Table 1.
+Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science.
+However, there is no marked difference in the performance of these models on ER tasks, we can say LLM has a weaker ability to complete ER in our task.
+This may be attributed to the limited contextual memory capabilities of the LLMs and this is we need to apply additional ER process.
+The normailzed Darwin shows the performance of our ER and normalization process, the result indicate that it not only achieve the ER task successfully, but also contribute to the NER and RE task.
+Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
 Model Task Precision Recall F1 score
@@
 ER 0.9127 0.8708 0.8913
-Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
-To gain a more comprehensive understanding of the improvements brought by each step in the standardization process, we conducted ablation experiments, the results are shown in Table 2 which clearly shows that each method contributes to the improvement of ER.
-The result indicates the most effective method is to optimize through the expert dictionary, which indicates that in addition to the accuracy of the expert dictionary, it can also fully cover all other labels except for the three labels representing the material.
-The improvement of ER-NF/A is also significant, as this process can eliminate most erroneous materials, and even has a stronger improvement on NER than ER-OL.
+To better understand the enhancements provided by each component in the ER/Normalization process, ablation studies were performed.
+The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER.
+The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels which are included in the dictionary.
+Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExactor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED.
+Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications.
+ER-N/F is also an indispensable part of the ER process.
+The normalized data is transformed into triples to construct the MKG, which comprises 162,605 nodes and 731,772 edges, as illustrated in the schematic diagram in Figure 5.
+Table 2: Result of the ablation experiment in normalization.
   Method NER F1 $\nabla$ RE F1 $\nabla$ ER F1 $\nabla$
@@
 ER-NF/A 84.01 -8.85 83.07 -5.20 84.54 -4.59
-ER-OL 88.59 -4.37 80.20 -8.07 50.00 -38.83
-Table 2: Result of the ablation experiment in normalization.
-Besides, ER-N/F pay attention to adjust the relation error between "Name" and "Formula", from a contribution perspective, it is also an indispensable part of ER.
-We also evaluated the LLM classifier by selecting 200 manually labeled data, and the statistical evaluation results in Table 3 shows the strong performance of LLM used in classification task.
-Model Accuracy F1 Score Specificity
-Darwin 0.9650 0.9581 0.9741
-Table 3: Binary classification performance through Fine-tuned Darwin.
-The triples are transferred from normalized data to construct the FMKG which contains 162,605 nodes and 731,772 edges, the schematic diagram shows in Figure 3.
-In the graph, there are 11 types of nodes and 13 types of relations.
-The FMKG store large amount of information in functional materials includes battery, catalyst and solar cell.
-For example, as the Figure 4 shows, by analyzing the frequency of material appearances in the battery domain, we observe that Co2O3 is the most frequently occurring material, followed by MoS2, graphite, TiO2, LiCoO2 etc which are commonly used in the battery domain.
-Likewise, it is evident that the most prevalent application within the battery field is lithium-ion batteries, which appear with a frequency significantly higher than other applications in the battery domain.
-These insights align with our understanding of these domains and materials, confirming that our knowledge graph effectively stores factual knowledge and offers a knowledge platform for material researchers.
-Figure 3: (a) Global schematic diagram of FMKG; (b) Local schematic diagram of FMKG.
-Figure 4: The image shows the top ten materials and applications in each field.
-For further validation of FMKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science.
+ER-ED 88.59 -4.37 80.20 -8.07 50.00 -38.83
+Figure 5: (a) Global schematic diagram of MKG; (b) Local schematic diagram of MKG.
+Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on publication year of paper.
+KG with earlier years applied similarity calculation algorithms to predict potential links, while KG with later years are used to verify how many sets of "Material Applications" in these predictions were reported in the following n years.
+The result shows in Figure 6.
+Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predication and the percentage of reported predication in next few year is displayed.
+The predictive capacity of the MKG is substantial, therefore, for each prediction, we only use the top 200 "material-application" pairs for validation.
+Looking at the results, as time progresses, the predicted materials are gradually verified.
+Using data from 2014, within nine years, 48.5% of the "material-application" predictions have been validated.
+This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields.
+The red, blue and green line are average percentage of reported link based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link predication, respectively.
+Network-based similarity has the highest accurate predication which indicates the effective of simulating the thinking of real material development using algorithms.
+The Figure 6 (b) is an example, visualizing the KG made from data before 2018, and the predicted "Material Application" reported within 5 years using network based algorithms.
+Figure 6: Validation of the graph completion. (a) Percentage of reported predication after years. (b) Example of predicated material - application using data before 2018.
+Conclusion and future work
+In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts.
+This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets.
+Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text.
+This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
+Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers.
+Our analysis confirms the effectiveness and credibility of the MKG.
+Furthermore, we have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding.
+This analysis has enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
+Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions.
+First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG.
+This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature.
+Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties.
+This enhancement significantly increases the granularity and practical utility of the KG.
+Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science.
+Lastly, the integration of our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset.
+This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
+Appendix
+A. Validation for MKG
+To validation of MKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science.
 After splitting these triples apart, annotator can obtain 1000 entities and 500 relations.
 The report from annotator shows in Table 3.
-“Name” is the highest priority core label and only serves as the head in the knowledge graph, so it will not be used as a relation.
-Notably, the labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation.
-This is due to the rigorous standardization of entities under these categories, using our expert dictionary.
+The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation.
+This is due to the rigorous normalization of entities under these categories, using our expert dictionary.
+However, strict normalization can also lead to certain entity loss, which is difficult to avoid.
 In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation.
@@
 A detailed analysis of "Name" and "Acronym" misclassifications reveals that most of these wrong entities originate from "Formula", which is due to the reasonable error of LLM"s binary classification of "Formula" and "Name" and ChemDataExtractor.
-The impact of this kind of error is not significant, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
+Fortunately, the impact of this kind of error is not significant and will not bring a significant influence, as fundamentally, the "Name", "Formula" and "Acronym" from same source all represent the same material.
+Table 3: Human evaluation metric on randomly selected triples.
 Label Number in total Entity disagree % disagree Relation disagree % disagree
@@
 Characterization 57 0 0 0 0
-Table 4: Human evaluation metric on randomly selected triples.
-Discussion
-In this research, we construct a functional material KG by fine-tuned LLMs and ensuring the traceability of all information.
-During the process, we encountered two main challenges and proposed corresponding solution: 1) Fine tuning LLM reduces the requirement for training set volume, but requires higher quality for each data in training set.
-when the quality of training dataset is limited, the performance of NER and RE is not good.
-Recursively generating training set data automatically from inference result is a good solution.
-High recall and precision data from each inference can be put into the training set after simple manual check.
-2) Fine-tuned LLMs perform well on NER and RE task, but for ER task, they can not shows a same.
-However, the ER is also important in graph construction.
-To address the issue, we apply several ER methods and technologies including ChemDataExtractor, mat2Vec, Word Embedding Clustering and construct an expert dictionary storing information about energy material.
-Theses methods and technologies can not only achieve high perfromance ER task, but also remove some erroneous entities and relations.
-To better demonstrate the reliability of our method and FMKG, we compare them with existing structural information extraction method and material KG.
-We employ a similar approach to NERRE task of John Dagdelen et al.
-[21] to extract structured information from the text.
-The difference is that we have applied ER method to correct entities and relations, and finish ER tasks before graph construction.
-We evaluated our pipeline on NERRE’s evaluation dataset.
-It worth to mention that cross dataset evaluation allow for flexibility of multiple plausible annotations of an entity under different labels (e.g., "cathode" in NERRE is placed in "Application" label while our expert tend to place it in "Descriptor" since "cathode" and "anode" are component of "battery" which is an application.), and presented the results in a Table 5.
-From the results, it is evident that our pipeline has slightly lower recall across labels compared to the NERRE task, but our method holds a competitive edge in precision.
-This is because we believe that the primary goal of KG construction is to ensure the accuracy of the data.
-Hence, in the process of ER, our strategy tends to sacrifice a portion of recall in exchange for higher precision.
-However, when considering both recall and precision, our method scores higher in F1 across the "Acronym", "Application", "Structure/Phase", and "Descriptor" labels compared to NERRE, with the improvement in "Acronym" being particularly significant due to its lower precision in NERRE.
-This underscores the advantage of our ER pipeline and indirectly attests to the credibility of the FMKG.
-Label Extraction recall (NERRE) Extraction precision (NERRE) Extraction F1 (NERRE) Extraction recall (DarwinN) Extraction precision (DarwinN) Extraction F1 (DarwinN)
-Formula 0.943 0.943 0.943* 0.828 1.000 0.906
-Name 0.692 1.000 0.818 0.667 1.000 0.800
-Acronym 0.667 0.400 0.500 0.571 1.000 0.727
-Descriptor 0.576 0.905 0.704 0.781 0.962 0.862
-Application 0.797 0.870 0.832 0.750 1.000 0.857
-Structure/Phase 0.754 0.920 0.829 0.613 0.905 0.730
-*Best F1 scores for each label.
-Table 5: Evaluation comparison of NERRE and DarwinN.
-Knowledge graphs specifically dedicated to general materials are rare.
-In our discussion, we draw a comparison with MatKG2 [15], showcasing the differences in construction methodologies between MatKG2 and FMKG, as depicted in the Figure 5.
-The construction of MatKG2 involves in a multi-step NER and RE pipeline, which can not catch the source of relations.
-In our method, we use end-to-end method that use a single LLM have been investigated for joint named entity recognition and relation extraction.
-It can not only catch keep source of each triple in KG so that enhance the factual basis of KG, but also perform better on RE task.
-Figure 5: Schematic comparison of FMKG and MatKG2.
-KGs are always dynamic, especially in cutting-edge field such as material science.
-The LLM internalizes a large amount of knowledge during the training process, it can perform good inference even when new nodes and relationships appear, making KG easy to update.
-Besides, our work is a starting point where other material researchers can continue build FMKG using our result as training dataset, using our dictionary, ChemDataExtractor and mat2vec to achieve ER task, our method can also be extend to full paper, extract more functional material information to construct a larger and more comprehensive functional material knowledge graph.
-Besides, our pipeline can also be used in other specialized domains to construct domain-specific KGs.
-Conclusion
-In this study, we introduce a new NLP pipeline for KG construction, which aim to efficiently extract the triples from unstructured scientific texts.
-The main feature of the method is that it can fine-tune the LLMs by annotating a small amount of data, and use the fine-tuned LLM to extract structured information from a large amount of unstructured text.
-The entire process does not rely on any prediction, which can maximize the authenticity and traceability of structured information.
-By employing this method, we construct a Functional Material Knowledge Graph (FMKG) contains the materials and their related knowledge from abstract of 150,000 peer-reviewed paper.
-After analyzing, we have demonstrated the effectiveness and credibility of FMKG.
-In addition, our method and KG have great potential in different dimensions.
-Firstly, enhancing the depth of structured information extraction to encompass entire research papers promises a richer, more detailed knowledge graph.
-This involves not only expanding the scope of data analyzed but refining the process to capture nuances within complex scientific texts.
-Secondly, refining entity labels within our system allows for a more precise categorization of data, including incorporating detailed attributes such as synthesis conditions or property parameters, which significantly improves the granularity and utility of the knowledge graph.
-Thirdly, the versatility of our NLP pipeline suggests its applicability across different scientific domains, offering a template for constructing domain-specific knowledge graphs beyond material science.
-Lastly, integrating FMKG with existing knowledge graphs like MatKG opens avenues for creating a more interconnected and comprehensive dataset, facilitating advanced research and application development in material science and beyond.
+B. Predication result
+As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023.
+It is important to note that these materials-application pairs were not ranked within the top five before 2017, which is why they do not appear in the figure for those earlier years.
+Stars mark the years these materials were featured in literature for their respective applications.
+The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas.
+This predication serves as an inspiration to materials researchers, accelerating progress in material development.
+For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER).
+This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
+C. Paramaters for LLM fine tuning
+epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0.
+warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: "LlamaDecoderLayer" token max length: 1024
+Figure 7: High rank predicated materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, the stars represent the material has been reported in this application.
+D. Experiments compute resources
+The compute resources were primarily consumed in LLM fine-tuning, involving the use of 8 GPUs for 9.02 hours with a 36% utilization rate, consuming 25.98 service units and using 14245 MiB of storage from a quota of 81920 MiB.

### v3

@@
 Introduction
-In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions.
+In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions [1].
 These conventional research approaches significantly impede the life-cycle of high-performance material research.
@@
 A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
-Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information of material science from the vast ocean of literature remains demanding.
-To obtain clearer sense of materials properties some structured database project such as Materials Project [2], OQMD [3], and NOMAD [4] were developed.
-However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [5].
-While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations.
-Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
+Despite the availability of comprehensive databases such as Scopus, Web of Science, and Crossref that offer structured search capabilities for scientific literature, extracting actionable materials science information from the extensive volumes of published research remains a daunting task.
+In response, databases like the Materials Project [3], OQMD [4], and NOMAD [5] have been developed to provide a clearer understanding of materials properties.
+While traditional databases are effective in storing structured data and handling basic queries, they may struggle to capture complex relations and infer new knowledge from the data [6].
+Specifically, traditional databases store data in tables with a predefined schema, organizing data into rows and columns, and establish relations between entities using primary and secondary keys.
+This architecture often requires complex and computationally intensive queries to extract interconnected data, which can hinder the effective mining of knowledge from various scientific findings.
 Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide.
-The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [6], [7].
-These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [8].
+The adoption of knowledge graphs offers several advantages, including enhanced data interoperability and visualization [7], the ability to infer new knowledge through relational data analysis or graph neural network [8], and improved data quality and consistency through structured representation [9], [10].
+These features make knowledge graphs particularly valuable for integrating diverse information sources, providing a unified view of a domain’s knowledge and more comprehensive contextual relationships, thereby facilitating more informed decision-making and discovery [11].
 However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [12].
@@
 In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science.
-Our contributions are highlighted in three key areas: 1) We propose a method to achieve named entity recognition (NER), relation extraction (RE) and entity resolution (ER) with high accuracy.
-Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet.
-This method also makes updating KG very convenient.
-2) We constructed the first accurate knowledge graph dedicated for materials, where researchers can easily get the information about material through query the MKG.
-3) We use a well-defined label system so our KG can be easily scale-up and potentially combined with other structured databases or KGs.
-4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
+Our contributions are as follows: (1) We propose a method to achieve high accuracy in Named Entity Recognition (NER), Relation Extraction (RE), and Entity Resolution (ER).
+This method allows for efficient conversion of unstructured text from complex research papers into structured information, effectively capturing crucial information and accelerating the research process in the early stages.
+This process not only retains the source information for each triplet but also makes updating the KG very convenient.
+(2) We constructed a dynamic MKG that supports rapid and automated updates, which is invaluable for the aggregation and processing of the growing global data on materials science.
+This capability ensures that researchers have access to the most current data, enhancing the predictive accuracy as the database expands.
+(3) We use a well-defined label system that enables easy scaling of our KG and potential integration with other structured databases or KGs like MatKG series and BoschKS [25].
+This well-defined labeling not only integrates professionalism into the construction of the knowledge graph but also acts as an effective link between researchers from both academia and industry.
+(4) We demonstrate a similarity calculation method based on the Jaccard Similarity for materials and applications.
+This protocol could enable the integration of knowledge and the discovery of novel materials for material applications from the wealth of scientific literature, which can facilitate the design process and inspire innovative ideas for future studies of material science researchers.
 Methods
@@
 Figure 4: Schematic comparison of MKG and MatKG2.
-To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [15], highlighting the differences in their construction methodologies as illustrated in Figure 4.
-MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation.
+To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG series [2], [18], highlighting the differences in their construction methodologies as illustrated in Figure 4.
+However, the relations in MatKG is based on statistical means, which can not guarantee high accuracy of the relations, especially when the article contains multiple materials.
+Therefore, our main comparison object is MatKG2.
+MatKG2 is constructed using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation.
 In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously.
@@
 This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
+Moreover, this method can automatically extract high-precision knowledge form corpus, leading to dynamic KGs.
 Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers.
 Our analysis confirms the effectiveness and credibility of the MKG.
-Furthermore, we have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding.
-This analysis has enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
+We have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding, which enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
 Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions.
@@
 Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science.
-Lastly, the integration of our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset.
+Lastly, the integration of our Material Knowledge Graph with existing general material KGs, such as MatKG series and BoschKS, paves the way for the creation of a more interconnected and expansive dataset.
 This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.

### v5

@@
-Construction and Application of Materials Knowledge Graph in Multidisciplinary Materials Science via Large Language Model
+neurips_2024.bib
 Abstract
-Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges for efficient discovery and integration of new materials.
+Knowledge in materials science is widely dispersed across extensive scientific literature, posing significant challenges to the efficient discovery and integration of new materials.
 Traditional methods, often reliant on costly and time-consuming experimental approaches, further complicate rapid innovation.
 Addressing these challenges, the integration of artificial intelligence with materials science has opened avenues for accelerating the discovery process, though it also demands precise annotation, data extraction, and traceability of information.
-To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques, integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges.
+To tackle these issues, this article introduces the Materials Knowledge Graph (MKG), which utilizes advanced natural language processing techniques integrated with large language models to extract and systematically organize a decade’s worth of high-quality research into structured triples, contains 162,605 nodes and 731,772 edges.
 MKG categorizes information into comprehensive labels such as Name, Formula, and Application, structured around a meticulously designed ontology, thus enhancing data usability and integration.
@@
 This structured approach not only streamlines materials research but also lays the groundwork for more sophisticated science knowledge graphs.
-Introduction
-In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions [1].
+11footnotetext: Corresponding authors: shaozhou@greendynamics.com.au; tong@greendynamics.com.au; wenjie.zhang@unsw.edu.au22footnotetext: These authors contributed equally to this work.
+1 Introduction
+In the contemporary information era, despite notable advancements, the creation and advancement of novel materials still heavily rely on traditional, time-consuming trial-and-error methods intertwined with chemical and physical intuitions.
 These conventional research approaches significantly impede the life-cycle of high-performance material research.
@@
 Yet, the electronic properties of materials across these different domains are highly related, and researchers in different domains can potentially learn from each other.
-To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [2].
+To accelerate the progress of materials research, there is a pressing need to efficiently integrate knowledge from various disciplines [venugopal2024matkg].
 However, this vital knowledge is scattered across a vast array of over 10 million scientific papers, covering diverse topics and disciplines such as materials preparation and functionalization methods, advanced materials characterization techniques, and the exploration of physical, chemical, and biological properties, along with their applications in fields like electronic devices, clean energy storage and transfer, and mechanical engineering.
@@
 A critical gap in current research infrastructure is the lack of an effective materials science database that can consolidate this scattered knowledge, facilitating easier access and interdisciplinary integration.
-Despite the availability of comprehensive databases such as Scopus, Web of Science, and Crossref that offer structured search capabilities for scientific literature, extracting actionable materials science information from the extensive volumes of published research remains a daunting task.
-In response, databases like the Materials Project [3], OQMD [4], and NOMAD [5] have been developed to provide a clearer understanding of materials properties.
-While traditional databases are effective in storing structured data and handling basic queries, they may struggle to capture complex relations and infer new knowledge from the data [6].
-Specifically, traditional databases store data in tables with a predefined schema, organizing data into rows and columns, and establish relations between entities using primary and secondary keys.
-This architecture often requires complex and computationally intensive queries to extract interconnected data, which can hinder the effective mining of knowledge from various scientific findings.
+Despite the existence of current databases of scientific literature such as Scopus, Web of Science, and Crossref, which offer ways to search for research papers based on specific labels, extracting useful information about material science from the vast ocean of literature remains demanding.
+To obtain a clearer sense of materials properties, some structured database projects such as Materials Project [jain2013materials], OQMD [saal2013materials], and NOMAD [draxl2019nomad] were developed.
+However, these databases contain many computational results obtained through techniques like Density Functional Theory (DFT) or Molecular Dynamics (MD) simulations [mrdjenovich2020propnet].
+While these computational databases can provide valuable references for predicting and understanding certain materials systems, they often face discrepancies with experimental observations.
+Therefore, there is an urgent need within the field of materials science for a database grounded in experimental research and practical information.
 Knowledge graph (KG) is a structured representation of information that models the controlled vocabulary and ontological relations of a topical domain as nodes and edges, enabling complex queries and insights that traditional databases cannot easily provide.
-The adoption of knowledge graphs offers several advantages, including enhanced data interoperability and visualization [7], the ability to infer new knowledge through relational data analysis or graph neural network [8], and improved data quality and consistency through structured representation [9], [10].
-These features make knowledge graphs particularly valuable for integrating diverse information sources, providing a unified view of a domain’s knowledge and more comprehensive contextual relationships, thereby facilitating more informed decision-making and discovery [11].
-However, the construction of knowledge graphs in specific fields always requires the participation of a large number of expert [12].
-This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [13].
-With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [14].
-For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [15].
-Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [16].
-A user-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [17].
-Recently, a material knowledge graph, MatKG and MatKG2, containing information on material properties, structure, and applications, has been developed [2], [18].
+The adoption of knowledge graphs offers several advantages, including enhanced data interoperability, the ability to infer new knowledge through relational data analysis, and improved data quality and consistency through structured representation [Ji], [Zhang2021NeuralSA].
+These features make knowledge graphs particularly valuable for integrating diverse information sources and providing a unified view of a domain’s knowledge, thereby facilitating more informed decision-making and discovery [MitchellACM].
+However, the construction of knowledge graphs in specific fields always requires the participation of a large number of experts [Zhong].
+This labor-intensive process not only limits the scalability of KGs but also impacts their performance and timeliness [PanShirui].
+With the rapid development of natural language processing (NLP), methods for extracting information from unstructured text and constructing knowledge graphs have become more efficient and accurate [NLPpromptKG].
+For instance, in 2016, the Metallic Materials Knowledge Graph (MMKG) was developed to store materials information from various web data resources [zhang2017mmkg].
+Knowledge graphs tailored to lithium-ion battery cathodes have been constructed, aimed at identifying potential new materials candidates [nie2022automating].
+User-friendly databases focusing on specific material types, such as Metal-Organic Framework Knowledge Graphs (MOF-KG), have been developed [an2023knowledge].
+Recently, a material knowledge graph, MatKG, and MatKG2, containing information on material properties, structure, and applications, has been developed [venugopal2024matkg], [MatKG2].
 However, these material knowledge graphs face even greater challenges.
-Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [19].
-Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [20].
+Firstly, although advancements in NLP technology have reduced the dependency on experts to a certain extent, training data still requires extensive annotation to enhance model accuracy [Largeannotation].
+Secondly, the construction of these knowledge graphs often involves predicting relationships between nodes to form triples, which means the entities represented in the KG are not always based on real instances [sousa2023explainable].
 This can diminish the authenticity and credibility of the KG.
 Additionally, this approach makes updating the knowledge graph difficult, as each new node introduced necessitates predicting its relationship with every other node, complicating the maintenance of a dynamic and accurate knowledge graph, especially in advanced fields like material science.
-Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the extraction and credibility of structured information [21], [22].
-The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [23],[24].
-This means improving the results of NER and RE without requiring a large amount of labor becomes possible and was adopt in our research.
+Acknowledging these challenges, the emergence of LLMs like GPT and LLaMA represents a breakthrough, offering new solutions to enhance the zero-shot method [tan2024retrievalenhancedmutationmasteryaugmenting], extraction, and credibility of structured information [GPT], [touvron2023llama].
+The fine-tuning technique of LLMs can significantly enhance their performance in specific domain text tasks through training with fewer samples [xie2024creation],[ExtractSI].
+This means improving the results of Named Entity Recognition (NER) and Relation Extraction (RE) without requiring a large amount of labor becomes possible and was adopted in our research.
 In this paper, we have achieved significant advancements in the development of Materials Knowledge Graph (MKG), a pioneering graph database tailored for the field of materials science.
-Our contributions are as follows: (1) We propose a method to achieve high accuracy in Named Entity Recognition (NER), Relation Extraction (RE), and Entity Resolution (ER).
-This method allows for efficient conversion of unstructured text from complex research papers into structured information, effectively capturing crucial information and accelerating the research process in the early stages.
-This process not only retains the source information for each triplet but also makes updating the KG very convenient.
-(2) We constructed a dynamic MKG that supports rapid and automated updates, which is invaluable for the aggregation and processing of the growing global data on materials science.
-This capability ensures that researchers have access to the most current data, enhancing the predictive accuracy as the database expands.
-(3) We use a well-defined label system that enables easy scaling of our KG and potential integration with other structured databases or KGs like MatKG series and BoschKS [25].
-This well-defined labeling not only integrates professionalism into the construction of the knowledge graph but also acts as an effective link between researchers from both academia and industry.
-(4) We demonstrate a similarity calculation method based on the Jaccard Similarity for materials and applications.
-This protocol could enable the integration of knowledge and the discovery of novel materials for material applications from the wealth of scientific literature, which can facilitate the design process and inspire innovative ideas for future studies of material science researchers.
-Methods
-Figure 1(a) presents the elaborate workflow of our study.
-Initially, through Named Entity Recognition (NER) and Relation Extraction (RE) tasks, we extract structural information pertinent to catalysts, batteries, and solar cells.
-Subsequent Entity Resolution (ER) and data normalization steps enable the integration of information from these distinct fields into a unified knowledge graph.
-Specifically, as outlined in Figure 1(b), the process begins with the manual annotation and normalization of the initial training dataset, essential for fine-tuning Large Language Models (LLMs) tailored for NERRE tasks.
-The inference dataset is then methodically divided into ten batches, facilitating the iterative refinement process that follows.
-The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [26], mat2vec [27], and our expertly curated dictionary.
+Our contributions are highlighted in three key areas: 1) We propose a method to achieve NER, RE, and entity resolution (ER) with high accuracy.
+Through this method, we can easily convert unstructured text into triples and retain the source information of each triplet.
+This method also makes updating KG very convenient.
+2) We constructed the first accurate knowledge graph dedicated to materials, where researchers can easily get information about the material by querying the MKG.
+3) We use a well-defined label system so our KG can be easily scaled up and potentially combined with other structured databases or KGs.
+4) We demonstrate a similarity calculation method based on Jaccard Similarity for materials and applications.
+2 Methods
+Figure 1 presents the elaborated pipeline of our study.
+The KG construction part can be divided into there tasks - ontology design, knowledge extraction, and entity resolution (ER).
+For domain-specific KG, the design of ontology often relies on experts in the field.
+This work forms an effective ontology by defining and summarizing a small number of papers through LLM.
+The knowledge extraction task begins with the manual annotation and normalization of the initial training dataset, annotated training set is used to finetune the LLM for NERRE tasks.
+Simultaneously, the inference dataset is divided into ten batches, facilitating the iterative refinement process that follows.
+The ER tasks are conducted using advanced NLP technologies, including ChemDataExtractor [CDE], mat2vec [mat2vec], and our expertly curated dictionary.
+These steps enable the integration of information from these distinct fields into a unified knowledge graph and clean the extracted data.
 After ER, we selectively enhance the training dataset with high-quality results to improve the model’s performance in subsequent iterations.
@@
 This methodology provides critical insights and recommendations for researchers in the materials science domain.
-Figure 1: (a) Workflow and (b) pipeline of the fine-tuned LLM for knowledge graph tasks.
-Data preparation and schema design
+Figure 1: Pipeline of the fine-tuned LLM for knowledge graph tasks.
+2.1 Data preparation and schema design
 Material experts annotated nine distinct categories from the abstracts of 75 research papers, forming the training dataset for the Large Language Models (LLMs).
 The structure of the data reflects the structure of the KG.
-As illustrated in Figure 2, the "Material" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes.
-Specifically, the "Material" node comprises sub-nodes such as "Formula," "Name," and "Acronym" for identifying materials.
+This structure summarizes ten articles through LLM and extracts key attributes.
+As illustrated in Figure 2 (a), the "Formula", "Name" and "Acronym" node serves as the central hub within the graph, linking to nodes that encapsulate its nomenclature, composition, and various attributes.
+Among these core attributes, the priority of the attribute pointed to by the arrow decreases in sequence.
 The "Structure/Phase" node describes the material’s physical state, while the "Application" node denotes its practical uses, and the "Property" node details its inherent characteristics.
@@
 Furthermore, the "Application" node is extended to include "Property", "Descriptor", and "Domain" nodes, indicating the specific attributes and the broader context of the material’s application.
+Among them, "Domain" is unique to "Application" and represents the field to which the application belongs.
 To maintain data provenance and traceability, each node is linked to a "Digital Object Identifier (DOI)" node.
-This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations was derived, ensuring data integrity within the graph structure.
-Figure 2: This schema represents the ontology of MKG (a) All the nodes in MKG connect to all their sources - "DOI".
-(b) Relations between all labels.
-(c) Relations between core labels, the more arrows pointing to the label, the lower the priority it has.
+This allows for source verification where querying the intersections of "DOI" node connections at both ends of a relation can pinpoint the source article from which the relations were derived, ensuring data integrity within the graph structure.
+Figure 2 (b) is an example of MKG, from which it can be seen that the core of MKG is to capture the connection and structure between materials and their applications.
+Figure 2: This schematic represents the (a) MKG schema and (b) an example of path in MKG between the "Name" node "Copper Indium Gallium Selenide" and "Application" node "Thin Films".
 Given the inherent complexity of sentences and the variability in terminologies across abstracts, a normalization process was employed after the initial extraction.
-This normalization ensured uniform representation of entities with similar meanings.
-For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed", ensuring consistency and facilitating effective training of LLMs.
+This normalization ensured a uniform representation of entities with similar meanings.
+For example, terms such as "Lithium-Ion Battery" and "Li-ion batteries" were standardized to "lithium-ion battery", while phrases like "solution casting method", "solvent post-treatment method", and "solution-based deposition" were simplified to "solution-processed".
+All annotated entities underwent this normalization process, ensuring consistency and facilitating effective training of LLMs.
 The field of materials encompasses a wide range of areas.
 At this stage, our priority is to focus on energy materials.
-We download 150,000 abstracts of peer-reviewed research articles of energy material science includes battery, solar cell and catalyst from the Web of Science.
-Each abstract was stored in a JSON file format, structured as "DOI - text," facilitating seamless processing and analysis.
-LLMs training, evaluation and inference
-The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [28], and Darwin [29].
-Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM for infer subsequent batches of data.
+We downloaded 150,000 abstracts of peer-reviewed research articles on energy material science, including batteries, solar cells, and catalysts, from the Web of Science.
+Each abstract was stored in a JSON file format, structured as "DOI - text", facilitating seamless processing and analysis.
+2.2 LLMs training, evaluation and inference
+The training dataset, composed of compiled data, was employed to fine-tune models including LLaMA 7b, LLaMA2 7b [touvron2023llama2], and Darwin [xie2023darwin].
+Upon obtaining high-quality results from the normalized inference, we iteratively retrain a fine-tuned LLM to infer subsequent batches of data.
 The models underwent training over 10 epochs with a batch size of 1.
 Additionally, 60 abstracts were annotated to assess the LLMs" performance.
-Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task"s ability to identify potential internal relations among relevant element sets.
+Our evaluation primarily focused on the NER capabilities of the model and preliminarily explored the RE task’s ability to identify potential internal relations among relevant element sets.
 Moreover, since we standardized entities during the data compilation phase, we also assessed the LLMs" effectiveness in standardizing the entities.
-Specifically, for NER, RE and ER task, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (fp) and false negatives (fn) to quantify the performance.
+Specifically, for NER, RE, and ER tasks, we employ a unified framework for evaluation based on precision, recall, and the F1 score, taking into account the instances of false positives (FP) and false negatives (fn) to quantify the performance.
 For NER, a true positive is a correctly identified entity, while for RE, it is a correctly identified relationship between entities, and for ER, it is a correctly standardized entity according to the schema.
@@
 After defining these terms, we can evaluate each task using standard precision, recall, and F1 score metrics.
-After evaluation, we choose a fine-tuned LLM, which demonstrates optimal performance in both NER and RE task, to iteratively infer the 150,000 abstracts.
-The output of inferences are organized in the "DOI - text - response" format.
-Consequently, the fine-tuned LLM not only extracts entities but also assigns them to appropriate labels, thereby accomplishing NER and RE tasks concurrently.
+After evaluation, we chose a fine-tuned LLM that demonstrates optimal performance in both NER and RE tasks to iteratively infer the 150,000 abstracts.
+The output of inferences is organized in the "DOI—text—response" format.
+Consequently, the fine-tuned LLM not only extracts entities but also assigns them appropriate labels, thereby accomplishing NER and RE tasks concurrently.
 Moreover, every entity and relation identified in the response is traceable, enhancing the integrity and utility of the data.
-Entity resolution
-The quality of KG is crucial for its credibility, to further check and correct the inference results before graph construction.
-To ensure the precision of these results, we initially employ ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts.
+2.3 Entity resolution
+The quality of KG is crucial for its credibility in checking and correcting the inference results before graph construction.
+To ensure the precision of these results, we initially employed ChemDataExtractor to identify chemical formulas and "Name-Acronym" pairs from abstracts.
 Subsequently, entities recognized by both the Large Language Model (LLM) and ChemDataExtractor are embedded using the mat2vec model.
 We analyze their similarities to rectify core entities and ensure accurate "Name" to "Acronym" associations.
-Given the frequent mislabeling of "Name" and "Formula", we progress to the Entity Resolution - Name/Formulas (ER-N/F) phase.
+Through this step, we can make the "Name" and "Formula" different from "Acronym" in MKG.
+Therefore, we have named this stage "ER-NF/A" ("Entity resolution - Name/Formula and Acronym").
+Given the frequent mislabeling of "Name" and "Formula", we progress to the "Entity resolution - Name and Formulas" ("ER-N/F") phase.
 Here, we refine the LLM using a specifically curated training set comprising 2,000 accurately labeled entities for binary classification, which is evaluated against an additional set of 200 labels.
-For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [30] algorithm to create the expert dictionary (ER-ED).
-This algorithm dynamically forms clusters based on vector similarity, without the need to predefine the number of clusters.
+For other labels, we implement a Density-Based Spatial Clustering of Applications with Noise [ester1996density] algorithm to create the "Entity resolution - expert dictionary" ("ER-ED").
+This algorithm dynamically forms clusters based on vector similarity without the need to predefine the number of clusters.
 Each cluster is named by material science experts, leading to the creation of an expert dictionary containing approximately 600 terms related to structures, phases, applications, and more.
@@
 To elevate the quality of the training dataset, we selectively integrate high-quality data from normalized inference outputs from each iteration into the training set, continuously monitoring and enhancing the performance of the fine-tuned LLM to ensure its efficacy.
-Knowledge graph construction
+2.4 Knowledge graph construction
 Given a collection of entities $E$ and relations $R$, a knowledge graph $\mathcal{K}=E\times R\times E$ is structured as a directed multi-relational graph.
@@
 To facilitate access to the detailed information, we also make the dataset available in the RDF and CSV format for straightforward data handling.
-Graph completion
-In the analysis depicted in Figure 3, we enhance the traditional Jaccard similarity, hereafter referred to as $J$, to form a comprehensive similarity metric for evaluating potential material-to-application links.
-The Jaccard similarity is mathematically defined as:
-$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$
-where $A$ and $B$ represent the attribute sets of two compared materials, such as their properties or applications.
-To refine our predication method, we extend this basic formulation by incorporating weighted factors that reflect the relevance and impact of each attribute, resulting in a modified similarity score for "Materials-Applications", $J_{\text{mat}}$.
-This score is computed as follows:
+2.5 Graph completion
+The process of Graph Completion (GC) is shown in Figure 3 (a), where we perform link prediction through GC, segmented into four primary stages: graph splitting, similarity calculation, validation and evaluation, and parameter optimization.
+This structured approach ensures a comprehensive exploration of the link prediction capabilities.
+Graph splitting is meticulously performed based on the chronological assignment of the nodes.
+Nodes encapsulated within a defined prediction window $\delta$ are utilized as $G_{\text{ver}}$, which are designated for the validation of predictions.
+Nodes preceding this predictive interval serve as $G_{\text{tra}}$, employed to train and refine the prediction models.
+During the similarity calculation stage, advanced graph algorithms and embeddings are applied on $G_{\text{tra}}$ to ascertain the similarity between the "Materials" and "Applications" nodes.
+This phase leverages an enhanced Jaccard similarity metric, partitioned into three distinct components: $S(m,a)$, $F(m,a)$, and $T(m,a)$, which collectively improve the specificity of predictions.
+Figure 3: (a)The process of MKG graph completion and (b) the schematic diagram of nodes comparison.
+In the validation and evaluation phase, the sorted "Material" and "Application" nodes undergo rigorous testing on $G_{\text{ver}}$.
+The effectiveness of the predictive model is assessed by tallying the incorrect predictions, represented as En, which serves as a crucial measure of performance.
+The final stage, parameter optimization, involves the meticulous adjustment of the parameters $S(m,a)$, $F(m,a)$, and $T(m,a)$ based on their performance on $G_{\text{ver}}$.
+This iterative process is crucial for minimizing the count of En and enhancing the overall accuracy of the GC method.
+This adaptive adjustment underscores the dynamic and responsive nature of the model in refining link prediction accuracy within the graph completion framework.
+It is worth mentioning that from a global perspective, we not only need to consider whether an "Application" may be similar to a "Material" node.
+We also need to consider whether the "Material" currently associated with this "Application" is similar to the potential "Material".
+Therefore, further improvement of Jaccard Similarity:
 $J_{\text{mat}}(A,B)=\sum_{i=1}^{n}w_{i}\cdot\frac{|A_{i}\cap B_{i}|}{|A_{i}\cup B_{i}|}$
 where $n$ is the number of attributes considered, $A_{i}$ and $B_{i}$ are the specific attributes of materials $A$ and $B$ respectively, and $w_{i}$ are the weights assigned to each attribute, signifying their importance in the context of material application.
-This approach mimics the investigative processes commonly employed by materials scientists when developing new materials.
+As shown in Figure 3 (b), this approach mimics the investigative processes commonly employed by materials scientists when developing new materials.
 In determining the suitability of new materials for specific applications, researchers systematically analyze their chemical and physical properties, as well as their structural characteristics, comparing these attributes to those of materials well-established in the target field.
-This enhanced method not only accounts for the quantity of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application.
+This enhanced method not only accounts for the number of overlapping attributes but also emphasizes the significance of each attribute in relation to the material’s potential application.
 A higher $J_{\text{modified}}$ score indicates a stronger likelihood of applicability between the material in question and the target application.
 Thus, materials scoring high on this metric are considered promising candidates for the specified applications.
-Figure 3: The process of network-based graph completion.
-In addition to the Enhanced Jaccard Similarity, we employ the TransE model [31].
+In addition to the Enhanced Jaccard Similarity, we employ the TransE model [bordes2013translating].
 TransE treats entities and relations as vectors in the same embedding space.
 The core idea of TransE is to model relations by interpreting them as translations in the embedding space.
-For a given relation $r$ between head $h$ and tail $t$, the relationship is said to hold if:
-$\|\mathbf{h}+\mathbf{r}-\mathbf{t}\|\approx 0$
-[…]
-Result
-Figure 4: Schematic comparison of MKG and MatKG2.
-To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG series [2], [18], highlighting the differences in their construction methodologies as illustrated in Figure 4.
-However, the relations in MatKG is based on statistical means, which can not guarantee high accuracy of the relations, especially when the article contains multiple materials.
-Therefore, our main comparison object is MatKG2.
-MatKG2 is constructed using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation.
+3 Result
+To illustrate the advantages of MKG more clearly, we conducted a comparative analysis with MatKG2 [MatKG2], highlighting the differences in their construction methodologies as illustrated in Figure 4.
+MatKG2 is built using a multi-step process for named entity recognition (NER) and relation extraction (RE) that does not retain the origin of each relation.
 In contrast, our approach employs an end-to-finish methodology utilizing a single large language model (LLM) designed to handle both NER and RE simultaneously.
 This not only preserves the source of each triple within the knowledge graph, enhancing its factual accuracy, but also demonstrates superior performance in the RE task.
-To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in the Table 1.
+Figure 4: Schematic comparison of MKG and MatKG2.
+To demonstrate the effectiveness of this pipeline, we evaluated the performance of each LLM on each task, as shown in Table 1.
 Darwin significantly outperforms both the LLaMA 7b and LLaMA2 7b models in the F1 scores for NER and RE tasks, suggesting that Darwin yields more effective results in text-related tasks in materials science.
-However, there is no marked difference in the performance of these models on ER tasks, we can say LLM has a weaker ability to complete ER in our task.
-This may be attributed to the limited contextual memory capabilities of the LLMs and this is we need to apply additional ER process.
-The normailzed Darwin shows the performance of our ER and normalization process, the result indicate that it not only achieve the ER task successfully, but also contribute to the NER and RE task.
-Table 1: Result of NER, RE and ER through Fine-tuned LLMs.
+However, there is no marked difference in the performance of these models on ER tasks; we can say LLM has a weaker ability to complete ER in our task.
+This may be attributed to the limited contextual memory capabilities of the LLMs, and this is why we need to apply additional ER processes.
+The normailzed Darwin shows the performance of our ER and normalization process, the result indicates that it not only achieves the ER task successfully, but also contributes to the NER and RE task.
+Table 1: Result of NER, RE, and ER through Fine-tuned LLMs.
 Model Task Precision Recall F1 score
+NER 0.1196 0.6869 0.2036
+MatBERT RE 0.0250 0.5696 0.0479
+ER 0.0928 0.5303 0.1579
 NER 0.6101 0.6216 0.6158
@@
 NER 0.9520 0.9083 0.9296
-Darwin (Normalized) RE 0.9039 0.8625 0.8827
+Darwin (ER) RE 0.9039 0.8625 0.8827
 ER 0.9127 0.8708 0.8913
@@
 The outcomes, detailed in Table 2, demonstrate that each technique positively affects the effectiveness of ER.
-The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels which are included in the dictionary.
-Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExactor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED.
+The results highlight that utilizing the expert dictionary (ER-ED) is the most beneficial method, enhancing accuracy comprehensively across nearly all labels that are included in the dictionary.
+Additionally, the improvements seen with ER-NF/A are noteworthy; this step is based on the ChemDataExtractor, effectively removes the majority of incorrect material identifications in core labels and offers even more substantial benefits to NER than to ER-ED.
 Furthermore, from the results of the ablation experiment, ER-N/F appears that its contribution is not as significant as the other two components, but considering the confusion between the material Formula and Name may have an impact on subsequent MKG applications.
@@
 Figure 5: (a) Global schematic diagram of MKG; (b) Local schematic diagram of MKG.
-Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on publication year of paper.
-KG with earlier years applied similarity calculation algorithms to predict potential links, while KG with later years are used to verify how many sets of "Material Applications" in these predictions were reported in the following n years.
-The result shows in Figure 6.
-Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predication and the percentage of reported predication in next few year is displayed.
-The predictive capacity of the MKG is substantial, therefore, for each prediction, we only use the top 200 "material-application" pairs for validation.
+Finally, to demonstrate the reliability and accuracy of our similarity algorithm, we divided MKG into two knowledge graphs based on the paper’s publication year.
+KG, with earlier years, applied similarity calculation algorithms to predict potential links, while KG, with later years, is used to verify how many sets of "Material-Applications" in these predictions were reported in the following n years.
+The result is shown in Figure 6.
+Specifically, in Figure 6 (a), the grey line uses only abstracts published before the year to make predictions, and the percentage of reported predictions in the next few years is displayed.
+The predictive capacity of the MKG is substantial; therefore, for each prediction, we only use the top 200 "material-application" pairs for validation.
 Looking at the results, as time progresses, the predicted materials are gradually verified.
@@
 This outcome demonstrates the effectiveness and reliability of the MKG, and it also lays the foundation for proposing new materials in some fields.
-The red, blue and green line are average percentage of reported link based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link predication, respectively.
-Network-based similarity has the highest accurate predication which indicates the effective of simulating the thinking of real material development using algorithms.
-The Figure 6 (b) is an example, visualizing the KG made from data before 2018, and the predicted "Material Application" reported within 5 years using network based algorithms.
-Figure 6: Validation of the graph completion. (a) Percentage of reported predication after years. (b) Example of predicated material - application using data before 2018.
-Conclusion and future work
-In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction aimed at efficiently extracting triples from unstructured scientific texts.
-This methodology distinguishes itself by enabling the fine-tuning of Large Language Models (LLMs) with minimally annotated datasets.
-Subsequently, the fine-tuned LLMs are utilized to extract structured information from extensive corpora of unstructured text.
-This entire process eschews predictive modeling, thereby enhancing the authenticity and traceability of the structured data extracted.
-Moreover, this method can automatically extract high-precision knowledge form corpus, leading to dynamic KGs.
-Utilizing this approach, we have constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers.
-Our analysis confirms the effectiveness and credibility of the MKG.
-We have explored potential connections between materials and applications through similarity analyses based on node similarity and graph embedding, which enabled us to validate inferred relations reliably, offering actionable insights for materials science research and fostering the field’s advancement.
-Additionally, our methodology and the resulting Knowledge Graph (KG) possess significant potential across various dimensions.
-First, by extending the depth of structured information extraction to cover entire research papers, we anticipate the development of a richer and more detailed KG.
-This extension involves not merely enlarging the dataset but also enhancing our extraction techniques to accurately capture subtle distinctions within complex scientific literature.
-Second, improving the precision of entity labeling in our pipeline enables more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties.
+The red, blue, and green lines are the average percentage of reported links based on network similarity, Jaccard similarity directly between "Formula" and "Application", and TransE for link prediction, respectively.
+Network-based similarity has the highest accurate prediction, which indicates the effectiveness of simulating the thinking of real material development using algorithms.
+The Figure 6 (b) is an example, visualizing the KG made from data before 2018 and the predicted "Material Application" reported within 5 years using network-based algorithms.
+Figure 6: Validation of the graph completion. (a) Percentage of reported prediction after years. (b) Example of predicted material - application using data before 2018.
+4 Conclusion, Discussion, and Future Work
+In this paper, we present a novel natural language processing (NLP) pipeline for Knowledge Graph (KG) construction, designed to efficiently extract triples from unstructured scientific texts.
+This methodology uniquely enables the fine-tuning of Large Language Models (LLMs) using minimally annotated datasets.
+The fine-tuned LLMs are then utilized to extract structured information from extensive corpora of unstructured text, bypassing predictive modeling to enhance the authenticity and traceability of the structured data extracted.
+Utilizing this approach, we constructed a Material Knowledge Graph (MKG) that encapsulates relations between materials and their associated entities, such as properties and applications, derived from the abstracts of 150,000 peer-reviewed papers.
+Our analysis not only confirms the effectiveness and credibility of the MKG but also enriches our understanding of the material science domain.
+Additionally, our methodology and the resulting KG hold significant potential across various dimensions:
+  1. 1.
+    Currently, our original text consists of article abstracts, which may omit a substantial amount of experimental methods and data.
+Fortunately, our method can be directly applied to full-text extraction, allowing us to more accurately capture subtle distinctions within complex scientific literature.
+  2. 2.
+    The ontology design of MKG has improved the precision of entity labeling in our pipeline and facilitates more accurate categorization of data, including the incorporation of intricate attributes such as synthesis conditions or material properties.
 This enhancement significantly increases the granularity and practical utility of the KG.
-Third, the adaptability of our NLP pipeline suggests its applicability to a broad range of scientific fields, providing a foundational template for constructing domain-specific KGs beyond the realm of materials science.
-Lastly, the integration of our Material Knowledge Graph with existing general material KGs, such as MatKG series and BoschKS, paves the way for the creation of a more interconnected and expansive dataset.
+  3. 3.
+    The adaptability of our NLP pipeline to a broad range of scientific fields suggests its potential as a foundational template for constructing domain-specific KGs beyond materials science.
+  4. 4.
+    Integrating our Material Knowledge Graph with existing KGs, such as MatKG, paves the way for the creation of a more interconnected and expansive dataset.
 This synergy facilitates not only advanced research but also the development of innovative applications in materials science and related domains.
-Appendix
-A. Validation for MKG
-To validation of MKG, we extract the 500 triples are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science.
-After splitting these triples apart, annotator can obtain 1000 entities and 500 relations.
-The report from annotator shows in Table 3.
+For future work, given the apparent predictive capacity of our method for link prediction, we plan to focus on:
+  1. 1.
+    Analyzing Historical Patterns: By incorporating additional attributes such as authorship, publication year, and affiliations into our Knowledge Graph, we focus on analyzing historical "social" patterns.
+For example, investigating whether specific groups or institutions consistently engage in the repurposing of materials.
+This analysis will allow us to understand how the dynamics of material repurposing are influenced by collaborative networks or institutional trends.
+We will assess which materials or types of materials are more frequently repurposed and explore the potential social dynamics that drive these trends.
+This approach aims to provide insights into the lifecycle and innovation patterns of materials based on past activities and collaborations.
+  2. 2.
+    Identifying Future Materials and Applications: After constructing a complete MKG, in-depth graph neural networks and algorithms can be developed and used to explore potential relations in the graph.
+By exploring the data within our enriched MKG, we will identify new potential applications for the most promising materials or explore the future trend of materials.
+This proactive approach seeks to unlock new uses for materials before they become mainstream, providing a competitive edge in materials science.
+Additionally, by incorporating time dynamics, we aim to enhance the utility of our pipeline, allowing for time-aware graph embedding techniques for graph completion.
+This will enable us to explore how material properties and applications evolve over time, providing a dynamic perspective on the development and use of materials in various fields.
+  3. 3.
+    Studying Cluster Formation: We plan to analyze the formation of clusters within the KG, which are based on the links between materials.
+Understanding these clusters will help reveal how materials are interconnected, which can aid in discovering the connections between different materials that were not apparent from isolated studies.
+5 Acknowledgments
+This study was undertaken with the assistance of resources and services from the GreenDynamics and National Computational Infrastructure (NCl), which is supported by the Australian Government, the department of education, and through the UNSW-NCl partner trial scheme.
+Microsoft Research also provides resources and support to this study through the Acceleration Foundational Models Research Grant.
+We thank UNSW Library for assisting our team in obtaining text and data mining agreements with publishers.
+[…]
+Appendix A Appendix
+A.1 Validation for MKG
+To validate MKG, we extract the 500 triples that are randomly extracted except the "DOI" and "Domain" nodes, which are informative information, and checked by the experts in relevant material science.
+After splitting these triples apart, the annotator can obtain 1000 entities and 500 relations.
+The report from the annotator is shown in Table 3.
 The labels "Application", "Structure/Phase", "Synthesis" and "Characterization" exhibit 100% accuracy in both entity and relation.
-This is due to the rigorous normalization of entities under these categories, using our expert dictionary.
+This is due to the rigorous normalization of entities under these categories using our expert dictionary.
 However, strict normalization can also lead to certain entity loss, which is difficult to avoid.
-In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, are applied a less stringent standardization process which lead to a certain degree of imprecision in entity and relation.
+In contrast, the "Descriptor" and "Property", characterized by their vast diversity and broad spectrum, apply a less stringent standardization process, which leads to a certain degree of imprecision in entity and relation.
 But the results are still satisfactory.
-"Name" and "Acronym" has also achieved a commendable 100% entity accuracy.
+"Name" and "Acronym" have also achieved a commendable 100% entity accuracy.
 However, when it comes to analyzing relations, the ChemDataExtractor encounters certain limitations.
@@
 Characterization 57 0 0 0 0
-B. Predication result
+A.2 prediction result
 As Figure 7 illustrates, we display the materials ranked in the top five for 2023 across various domains, alongside their rankings in previous years based on available data up to 2023.
@@
 The figure emphasizes the significance and emerging utility of these materials by highlighting the top five in applications such as catalysts, batteries, fuel/solar cells, and other fields from 2017 to 2023, indicating their growing importance and potential to revolutionize technology in these areas.
-This predication serves as an inspiration to materials researchers, accelerating progress in material development.
+This prediction serves as an inspiration to materials researchers, accelerating progress in material development.
 For example, consider the material NiFeN HC/NF; it not only excels in the Oxygen Evolution Reaction (OER) but has also demonstrated its capability as a catalyst for the Hydrogen Evolution Reaction (HER).
 This means that NiFeN HC/NF is a promising bifunctional electrocatalyst for water electrolysis, aligning with our understanding of its potential in dual-function applications.
-C. Paramaters for LLM fine tuning
+A.3 Paramaters for LLM fine tuning
 epochs: 10 train batch size (per device): 1 eval batch size (per device): 1 gradient accumulation steps: 2 evaluation strategy: "no" save strategy: "steps" save steps: 500 save total limit: 1 learning rate: 2e-5 weight decay: 0.
-warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: "LlamaDecoderLayer" token max length: 1024
-Figure 7: High rank predicated materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, the stars represent the material has been reported in this application.
-D. Experiments compute resources
-The compute resources were primarily consumed in LLM fine-tuning, involving the use of 8 GPUs for 9.02 hours with a 36% utilization rate, consuming 25.98 service units and using 14245 MiB of storage from a quota of 81920 MiB.
+warmup ratio: 0.03 lr scheduler type: "cosine" logging steps: 1 tf32: False fsdp: "full shard auto wrap" fsdp transformer layer cls to wrap: ’LlamaDecoderLayer’ token max length: 1024
+Figure 7: Top five potential materials in (a) Catalyst, (b) Fuel/Solar cell, (c) Battery and (d) Other fields, predicted on all available data (2014 - 2023) and their rank trends in the past years (2014 - related year).
+The stars represent the material has been reported in the related year.
+A.4 Experiments compute resources
+The compute resources were primarily consumed in LLM fine-tuning, which involved using 8 GPUs for 9.02 hours with a 36% utilization rate, 25.98 service units, and 14245 MiB of storage from a quota of 81920 MiB.